# Computation and Communication Co-scheduling for Timely Multi-Task Inference at the Wireless Edge

### v1

@@ Abstract
In multi-task remote inference systems, an intelligent receiver (e.g., command center) performs multiple inference tasks (e.g., target detection) using data features received from several remote sources (e.g., edge sensors). Key challenges to facilitating timely inference in these systems arise from (i) limited computational power of the sources to produce features from their inputs, and (ii) limited communication resources of the channels to carry simultaneous feature transmissions to the receiver. We develop a novel computation and communication co-scheduling methodology which determines feature generation and transmission scheduling to minimize inference errors subject to these resource constraints. Specifically, we formulate the co-scheduling problem as a weakly-coupled Markov decision process with Age of Information (AoI)-based timeliness gauging the inference errors. To overcome its PSPACE-hard complexity, we analyze a Lagrangian relaxation of the problem, which yields gain indices assessing the improvement in inference error for each potential feature generation-transmission scheduling action. Based on this, we develop a maximum gain first (MGF) policy which we show is asymptotically optimal for the original problem as the number of inference tasks increases. Experiments demonstrate that MGF obtains significant improvements over baseline policies for varying tasks, channels, and sources.
@@ I Introduction
The simultaneous advances in machine learning and communication technologies have spurred demand for intelligent networked systems across many domains [1, 2]. These systems, whether for commercial or military purposes, often rely on timely information delivery to a remote receiver for conducting several concurrent decision-making and control tasks [3]. For example, consider intelligence, surveillance, and reconnaissance (ISR) [4] objectives within military operations. A command center may employ signals transmitted from several dispersed military assets, e.g., unmanned aerial vehicles (UAVs), to simultaneously classify friendly versus hostile agents, track the positions of targets, and detect anomalous sensor data. Similarly, in intelligent transportation [5], near real-time prediction of road conditions, vehicle trajectories, and other tasks is crucial for traffic management and safety.
@@
As the number and complexity of learning tasks in such applications continues to rise, there are two salient challenges to facilitating timely multi-task remote inference (MTRI). First, there are limited wireless resources (e.g., orthogonal frequency channels) available for information transmission from sources to the receiver at the network edge. As a result, the sources may locally construct low-dimensional feature representations of their high-dimensional signal observations (e.g., video streams) to send in lieu of the raw measurements. This may be developed, for example, by splitting the neural network for each task at a designated cut layer, and implementing the two parts at the source and receiver, respectively [6]. However, this also leads to the second challenge: the sources, often edge devices, have heterogeneous on-board computational capabilities, limiting their ability to simultaneously construct multiple features required by different tasks. Due to these resource limitations, the features at the receiver may not always reflect the freshest source information.
@@
It is thus critical to ascertain which tasks require feature updates most urgently at any given time, i.e., to determine where to focus available MTRI resources. Age of Information (AoI), introduced in [7, 8], can provide a useful measure of information freshness of the receiver. Specifically, consider packets sent from a source to a receiver: if $U(t)$ is the generation time of the most recently received packet by time $t$, then the AoI at time $t$ is the difference between $t$ and $U(t)$. Recent works on remote inference [9, 3, 10, 11] have shown that the inference errors for different tasks can be expressed as functions of AoI, and that surprisingly, these functions are not always monotonic. Additionally, AoI can be readily tracked in an MTRI system on a per-task basis, making it a promising metric for determining how to prioritize resource allocation.
Motivated by this, we pose the following research question:
@@ I-A Outline and Summary of Contributions
  • Leveraging these gain indices, we propose a novel maximum gain first (MGF) policy to solve the original problem, iteratively scheduling features/tasks with maximum gain until capacity is reached (Sec. V-A). The MGF policy is a special case of the re-optimized fluid (ROF) policy introduced in [12] for general weakly coupled MDPs. We prove that in the MTRI problem, our MGF policy achieves asymptotic optimality at a rate of $O(\frac{1}{\sum_{m=1}^{M}\sqrt{rk_{m}}})$, where $rk_{m}$ is the number of inference tasks per source $m$ and $M$ is the total number of sources (Theorem 1, Sec. V-B). Notably, this optimality gap is tighter than the $O(\frac{1}{\sqrt{\sum_{m=1}^{M}rk_{m}}})$ bound established in [12]. Our scheduling results are applicable to any bounded penalty functions of AoI with multiple resource constraints.
  • We conduct numerical experiments to demonstrate our policy on both synthetic and real-world inference tasks (Sec. VI). In the latter case, we consider vehicular inference tasks (image segmentation and traffic prediction) where roadside sensors equipped with cameras are used as the MTRI sources. We find that MGF significantly outperforms baseline policies in terms of cumulative errors as the number of tasks, channels, and sources are varied. A widening margin is observed as the number of tasks increases, consistent with our optimality analysis.
@@ I-B Related Works
AoI has become a widely-used metric in the analysis and optimization of systems including communication networks [13, 14], control systems [15, 16], remote estimation [17, 18, 19], and remote inference [20, 9, 3]. Previous works [20, 9, 3] have demonstrated that the performance of remote inference systems depends on the AoI of the features they utilize; specifically, representing inference error as a function of AoI. In this paper, we consider the more challenging MTRI case with multiple information sources, an edge receiver, and multiple inference tasks for each source. Motivated by the prior work, we consider the dependency of inference error for each task on the AoI of features delivered to the receiver. Notably, the inference error function in our case can be monotonic or non-monotonic with AoI.
@@ II System Model
Figure 1: An MTRI system with $M$ sources, $N$ channels, $k_{m}$ feature generators per source $m$ (in this figure, $k_{m}=2$), and $K$ predictors.
@@ II-A Overview
We consider the multi-task remote inference (MTRI) system in Fig. 1. $M$ sources are connected to an intelligent receiver via $N$ wireless channels. For example, in an ISR system, UAVs equipped with cameras can act as sources, transmitting processed video frames to a central command center for further analysis. At every time slot $t$, each source $m$ observes a time-varying signal $X_{m,t}\in\mathcal{X}_{m}$, where $\mathcal{X}_{m}$ represents the set of possible observations, e.g., possible values of video frames captured by a camera. Sources will progressively generate low-dimensional feature representations of their observations for communication-efficient transmission over the $N$ wireless channels when they are scheduled.
At each time $t$, the receiver employs multiple predictors trained to infer targets based on received source features. Specifically, for each source $m$, the receiver aims to infer $k_{m}$ targets $(Y_{m,1,t},Y_{m,2,t},\ldots,Y_{m,k_{m},t})$. These targets can represent various inference tasks, e.g., object detection or segmentation, depending on the nature of the observations and the goals of the MTRI system. In the system, there are a total of $K=\sum_{m=1}^{M}k_{m}$ inference tasks. The tuple $(m,j)$ uniquely identifies the $j$-th inference task of source $m$.
@@ II-B Computation Model
Each source $m$ is equipped with $k_{m}$ pre-trained feature generators. The $j$-th feature generator of source $m$, designed for the $(m,j)$-th inference task, is denoted by a function $\phi_{m,j}:\mathcal{X}_{m}\mapsto\mathcal{Z}_{m,j}$. This function takes the observation $X_{m,t}\in\mathcal{X}_{m}$ as input and generates a feature $F_{m,j,t}=\phi_{m,j}(X_{m,t})\in\mathcal{Z}_{m,j}$, where $\mathcal{Z}_{m,j}$ is the set of possible features generated by $\phi_{m,j}(\cdot)$. To account for computational resource limitations, we assume it is not feasible to activate all feature generators at every time slot. Specifically, for source $m$, at most $C_{m}$ feature generators can be activated at any given time.
@@ II-C Communication Model
As illustrated in Fig. 1, $N$ wireless channels are shared among the $M$ sources. If scheduled at time $t$, the $(m,j)$-th feature generator produces $F_{m,j,t}$ and transmits to the receiver using $n_{m,j}$ channels. For simplicity, we assume perfect channels, i.e., features sent at time slot $t$ are delivered error-free at time slot $t+1$. However, our results can be extended to accommodate erasure channels, where data loss may occur.
Due to the limited number of channels, at any given time $t$, only features for a subset of inference tasks can be transmitted. Consequently, the receiver may not have fresh features for all tasks. If the most recently delivered feature for the $(m,j)$-th inference task was generated $\Delta_{m,j}(t)$ time slots ago, then the feature at the receiver is represented as
$F_{m,j,t-\Delta_{m,j}(t)}=\phi_{m,j}(X_{m,t-\Delta_{m,j}(t)}),$
where $\Delta_{m,j}(t)$ is its age of information (AoI) [8, 3]. Let $U_{m,j}(t)$ be the generation time of the most recent delivered feature. Then, the AoI can be formally defined as:
$\Delta_{m,j}(t)=t-U_{m,j}(t),$ (1)
@@ II-D Inference Model
where $P_{Y_{m,j,t},X_{m,t-\delta}}$ is the joint distribution of the target $Y_{m,j,t}$ and the observation $X_{m,t-\delta}$, and $L_{m,j}(y,\hat{y})$ is the loss function for the task that measures the loss incurred when the actual target is $y$ and the inference result is $\hat{y}$ (e.g., cross-entropy loss for a classification task).
@@ III-A Scheduling Policy and Optimization
$\bar{p}_{opt}=$
$\;\inf_{\pi\in\Pi}\sum_{t=0}^{\infty}\frac{\gamma^{t}}{K}\sum_{m=1}^{M}\sum_{j=1}^{k_{m}}\mathbb{E}_{\pi}\left[w_{m,j}p_{m,j}(\Delta_{m,j}(t))\right],$ (4)
$~{}\mathrm{s.t.}\sum_{j=1}^{k_{m}}\pi_{m,j}(t)\leq C_{m},t=0,1,\ldots,m=1,\ldots,M,$ (5)
$~{}~{}~{}~{}\sum_{m=1}^{M}\sum_{j=1}^{k_{m}}\pi_{m,j}(t)n_{m,j}\leq N,~{}~{}t=0,1,2,\ldots,$ (6)
where $w_{m,j}\geq 0$ is the weight (e.g., priority) associated with the $(m,j)$-th inference task, and the discount factor $10<\gamma<$ 1 quantifies the diminishing importance of an inference task over time. At most $C_{m}$ feature generators for source $m$ can compute features at time $t$. Transmitting features for the $(m,j)$-th inference task requires $n_{m,j}$ of the $N$ wireless channels available. For each task, its inference error, $p_{m,j}(\Delta_{m,j}(t))$, depends on its AoI $\Delta_{m,j}(t)$ at time slot $t$, indicating the freshness of the feature used for inference.
@@ III-B Weakly Coupled MDP Formulation
The problem (III-A)-(6) is a weakly coupled Markov decision process (MDP) [12, 30, 32] with $K$ sub-MDPs (referred to as arms in the bandit literature), one per inference task $(m,j)$ across sources $m$. The state of each $(m,j)$-th MDP at each time $t$ is represented by the AoI $\Delta_{m,j}(t)$. The action is $\pi_{m,j}(t)$, and its per-timeslot cost is $\gamma^{t}p_{m,j}(\Delta_{m,j}(t))$ with discount factor $10<\gamma<$ 1. We can see that the state evolution defined in (3) and the cost of each MDP depends only on its current state and action. However, the actions for all MDPs $(\pi_{m,j}(t))_{\forall m,j}$ need to satisfy the constraints in (5)-(6). This interdependence of actions across MDPs through multiple resource constraints, despite independent state transitions and costs, makes the overall problem (III-A)-(6) a weakly coupled MDP [12, 30, 32]. Weakly coupled MDPs are PSPACE-hard because the number of states and actions grow exponentially with the number of sub-MDPs.
The restless multi-armed bandit (RMAB) problem is a special case of the weakly coupled MDP in (III-A)-(6). RMAB considers a single resource constraint, whereas our problem involves multiple resource constraints (5)-(6). The PSPACE-hard complexity of RMABs can be overcome by using Whittle indices [28, 9, 3, 24, 21, 33, 26], gain indices [10, 19, 27], and linear programming-based indices [34] to construct asymptotically optimal policies, provided indexability and/or global attractor conditions are satisfied. However, these RMAB policies cannot be directly applied to our more general problem due to the presence of multiple resource constraints, which requires us to develop a new solution approach.
Solution approach. In Sec. IV&V, we follow the approach depicted in Fig. 2 to solve (III-A)-(6). We begin by deriving a relaxed Lagrangian problem. We then utilize the resulting solution to construct a maximum gain first (MGF) policy (Algorithm 1) for (III-A)-(6). Theorem 1 will demonstrate that the MGF policy becomes asymptotically optimal as the number of inference task $k_{m}$ for each source $m$ increases.

### v2

# Computation and Communication Co-scheduling for Multi-Task Remote Inference

@@ Abstract
In multi-task remote inference systems, an intelligent receiver (e.g., command center) performs multiple inference tasks (e.g., target detection) using data features received from several remote sources (e.g., edge devices). Key challenges to facilitating timely inference in these systems arise from (i) limited computational power of the sources to produce features from their inputs, and (ii) limited communication resources of the channels to carry simultaneous feature transmissions to the receiver. We develop a novel computation and communication co-scheduling methodology which determines feature generation and transmission scheduling to minimize inference errors subject to these resource constraints. Specifically, we formulate the co-scheduling problem as a weakly-coupled Markov decision process with Age of Information (AoI)-based timeliness gauging the inference errors. To overcome its PSPACE-hard complexity, we analyze a Lagrangian relaxation of the problem, which yields gain indices assessing the improvement in inference error for each potential feature generation-transmission scheduling action. Based on this, we develop a reoptimized maximum gain first (MGF) policy. We show that this policy is asymptotically optimal for the original problem as the number of inference tasks and the available communication and computation resources increase, provided the ratio among them remains fixed. Experiments demonstrate that reoptimized MGF obtains significant improvements over baseline policies for varying numbers of tasks, channels, and sources.
Index Terms:
Scheduling, resource allocation, age of information, multi-task inference, edge computing.
@@ I Introduction
The simultaneous advances in machine learning and communication technologies have spurred demand for intelligent networked systems across many domains [2, 3]. These systems, whether for commercial or military purposes, often rely on timely information delivery to a remote receiver for conducting several concurrent decision-making and control tasks [4]. For example, consider intelligence, surveillance, and reconnaissance (ISR) [5] objectives within military operations. A command center may employ signals transmitted from several dispersed military assets, e.g., unmanned aerial vehicles (UAVs), to simultaneously classify friendly versus hostile agents, track the positions of targets, and detect anomalous sensor data. Similarly, in intelligent transportation [6], near real-time prediction of road conditions, vehicle trajectories, and other tasks is crucial for traffic management and safety. Smart retail management system requires to infer current inventory and to classify customer reactions.
@@
As the number and complexity of learning tasks in such applications continues to rise, there are two salient challenges to facilitating timely multi-task remote inference (MTRI). First, there are limited wireless resources (e.g., orthogonal frequency channels) available for information transmission from sources to the receiver at the network edge. The observed information at the sources, often edge devices, can be high dimensional and transmitting the high dimensional data to cloud server requires a significant amount of communication resources. Due to limited communication resources, the delivered information may end up being stale. Because of advancement of hardware at the edge devices, questions arise as to whether this receiver (e.g., server)-only processing is efficient. Edge device computing has a potential to reduce communication resources needed. Instead of sending high dimensional signal values, AI-powered edge devices may locally construct low-dimensional feature representations of their high-dimensional signal observations (e.g., video streams) to send in lieu of the raw measurements. This may be developed, for example, by splitting the neural network for each task at a designated cut layer, and implementing the two parts at the edge device and receiver, respectively [7]. However, this also leads to the second challenge: the sources, often edge devices, have heterogeneous on-board computational capabilities, limiting their ability to simultaneously construct multiple features required by different tasks. For example, smart glasses may have low compute power, whereas high compute power can be installed on a vehicles. Hence, the feature computational ability of smart glasses is significantly lower compared to vehicles.
@@
Due to these resource limitations, the features at the receiver may not always reflect the freshest source information. It is thus critical to ascertain which tasks require feature updates most urgently at any given time, i.e., to determine where to focus available MTRI resources. Age of Information (AoI), introduced in [8, 9], can provide a useful measure of information freshness of the receiver. Specifically, consider packets sent from a source to a receiver: if $U(t)$ is the generation time of the most recently received packet by time $t$, then the AoI at time $t$ is the difference between $t$ and $U(t)$. Recent works on remote inference [10, 4, 11, 12] have shown that the inference errors for different tasks can be expressed as functions of AoI, and that surprisingly, these functions are not always monotonic. Additionally, AoI can be readily tracked in an MTRI system on a per-task basis, making it a promising metric for determining how to prioritize resource allocation. Motivated by this, we pose the following research question:
@@ I-A Outline and Summary of Contributions
  • Leveraging these gain indices, we propose a novel maximum gain first (MGF) policy (Algorithm 1) to solve the original problem, iteratively scheduling features/tasks with maximum gain until capacity is reached (Sec. V-A). The MGF policy is a special case of the re-optimized fluid (ROF) policy introduced in [13] for general weakly coupled MDPs. We prove that in the MTRI problem, our MGF policy achieves asymptotic optimality at a rate of $O(\frac{1}{\sum_{m=1}^{M}\sqrt{rk_{m}}})$, where $rk_{m}$ is the number of inference tasks per source $m$ and $M$ is the total number of sources (Theorem 1, Sec. V-B). Notably, this optimality gap is tighter than the $O(\frac{1}{\sqrt{\sum_{m=1}^{M}rk_{m}}})$ bound established in [13]. Our scheduling results are applicable to any bounded penalty functions of AoI with multiple resource constraints. We also provide Algorithm 2 by reducing the number of optimization variables of Algorithm 1.
  • We conduct numerical experiments to demonstrate our policy on synthetic and real-world inference tasks (Sec. VI). For synthetic evaluations, we use three different types of inference error functions that are widely used in AoI literature [14, 15]. For real-world inference tasks, we consider remote robot car detection and vehicular inference tasks (image segmentation and traffic prediction). In the remote robot car detection experiment, $4$ sources and $5$ robot cars are used, where source 1 observes 2 of the cars, while the remaining 3 sources each monitor a single car. In the vehicular inference tasks, roadside sensors equipped with cameras are used as the MTRI sources. We find that MGF significantly outperforms baseline policies in terms of cumulative errors as the number of tasks, channels, and sources are varied. A widening margin is observed as the number of tasks increases, consistent with our optimality analysis.
@@ I-B Related Works
The concept of Age of Information (AoI) has attracted significant research efforts; see, e.g., [16, 9, 17, 18, 19, 20, 21, 22, 15, 23, 24, 25, 26, 14, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37] and a survey [38]. Initially, research efforts were centered on analyzing and optimizing linear functions of AoI, as a performance metric of communication networks [9, 17, 18, 19, 20]. Recently, researchers have shifted their efforts towards optimizing the performance of real-time applications, such as remote estimation [39, 26, 36, 26, 32], remote inference [16, 10, 4, 40], and control systems [27, 37], by leveraging AoI as a tool. Previous works [16, 10, 4] have demonstrated that the performance of remote inference systems depends on the AoI of the features they utilize; specifically, representing inference error as a function of AoI. In this paper, we consider the more challenging MTRI case with multiple information sources, an edge receiver, and multiple inference tasks for each source. Motivated by the prior work, we consider the dependency of inference error for each task on the AoI of features delivered to the receiver. Notably, the inference error function in our case can be monotonic or non-monotonic with AoI. This paper is also related to the field of signal-agnostic remote estimation. The prior studies [15, 26, 32, 41, 39, 27, 33, 14] in signal-agnostic remote estimation focused on Gaussian and Markovian processes and found that the estimation error can be represented as a function of AoI values.
@@
(a) An MTRI system.
(b) $m$-th source with $k_{m}$ inference tasks.
Figure 1: System Model: (a) The multi-task remote inference (MTRI) system, where $M$ sources are connected to an intelligent receiver via $N$ wireless channels. Each source $m$ observes the high-dimensional signal $X_{m,t}$. At each time slot $t$, the goal of the receiver is to predict time-varying targets $(Y_{1,t},Y_{2,t},\ldots,Y_{M,t})$, where the target $Y_{m,t}$ is a tuple of $k_{m}$ time-varying signals $(Y_{m,1,t},Y_{m,2,t},\ldots,Y_{m,k_{m},t})$ of source $m$. (b) To address the $k_{m}$ inference tasks, each source $m$ generates and sends $k_{m}$ features, each associated with a single inference task. The receiver uses $k_{m}$ predictors to perform the $k_{m}$ inference tasks of source $m$, where $j$-th predictor for source $m$ predicts the target $Y_{m,j,t}$ using the most recently delivered feature from the $j$-th feature generator of source $m$.
@@ II-A Overview
We consider the multi-task remote inference (MTRI) system, as illustrated in Fig. 1(a), where $M$ sources are connected to an intelligent receiver via $N$ wireless channels and the receiver. Each source $m$ can be equipped with one sensor or multiple sensors addressing multiple tasks as depicted in Fig. 1(b); for example, in an ISR system, an UAV equipped with a camera can act as a source, transmitting processed video frames to a central command center for multiple tasks such as object recognition and anomaly detection. Another example is a robot using a LIDAR sensor for 3D mapping and a chemical sensor for hazardous material detection. At each time slot $t$, each source $m$ observes a high dimensional time-varying signal $X_{m,t}\in\mathcal{X}_{m}$, where $\mathcal{X}_{m}$ represents the set of possible observations, e.g., possible values of video frames captured by a camera. Sources will progressively generate low-dimensional feature representations of their observations for communication-efficient transmission over the $N$ wireless channels when they are scheduled.
At each time $t$, the receiver employs multiple predictors trained to infer targets based on received source features. Specifically, for each source $m$, the receiver aims to infer $Y_{m,t}$ which is a tuple of $k_{m}$ time-varying targets $(Y_{m,1,t},Y_{m,2,t},\ldots,Y_{m,k_{m},t})$ of source $m$. These targets can represent various inference tasks, e.g., object detection, segmentation, anomaly detection, 3D mapping, depending on the nature of the observations and the goals of the MTRI system. In the system, there are a total of $K=\sum_{m=1}^{M}k_{m}$ inference tasks. The tuple $(m,j)$ uniquely identifies the $j$-th inference task of source $m$.
@@ II-B Computation Model
Each source $m$ is equipped with $k_{m}$ pre-trained feature generators. The $j$-th feature generator of source $m$, designed for the $(m,j)$-th inference task, is denoted by a function $\phi_{m,j}:\mathcal{X}_{m}\mapsto\mathcal{Z}_{m,j}$. This function takes the observation $X_{m,t}\in\mathcal{X}_{m}$ as input and generates a feature $\phi_{m,j}(X_{m,t})\in\mathcal{Z}_{m,j}$, where $\mathcal{Z}_{m,j}$ is the set of possible features generated by $\phi_{m,j}(\cdot)$. To account for computational resource limitations, we assume it is not feasible to activate all feature generators at every time slot. Specifically, for source $m$, at most $C_{m}$ feature generators can be activated at any given time.
@@ II-C Communication Model
As illustrated in Fig. 1(a), $N$ wireless channels are shared among the $M$ sources. If scheduled at time $t$, the $(m,j)$-th feature generator produces $\phi_{m,j}(X_{m,t})$ and transmits to the receiver using $n_{m,j}$ channels. For simplicity, we assume perfect channels, i.e., features sent at time slot $t$ are delivered error-free at time slot $t+1$. However, our results can be extended to accommodate erasure channels, where data loss may occur.
Due to the limited number of channels, at any given time $t$, only features for a subset of inference tasks can be transmitted. Consequently, the receiver may not have fresh features for all tasks. If the most recently delivered feature for the $(m,j)$-th inference task was generated $\Delta_{m,j}(t)$ time slots ago, then the feature at the receiver is $\phi_{m,j}(X_{m,t-\Delta_{m,j}(t)}),$ where $\Delta_{m,j}(t)$ is its age of information (AoI) [9, 4]. Let $U_{m,j}(t)$ be the generation time of the most recent delivered feature. Then, the AoI can be formally defined as:
$\Delta_{m,j}(t):=t-U_{m,j}(t),$ (1)
@@ II-D Inference Model
[…]
@@ III-A Scheduling Policy and Optimization
$\bar{p}_{opt}$ $=\inf_{\pi\in\Pi}\sum_{t=0}^{\infty}\frac{\gamma^{t}}{K}\sum_{m=1}^{M}\sum_{j=1}^{k_{m}}\mathbb{E}_{\pi}\left[w_{m,j}p_{m,j}(\Delta_{m,j}(t))\right],$ (4)
$\quad~~\mathrm{s.t.}{\sum_{j=1}^{k_{m}}\pi_{m,j}(t)\leq C_{m},}t=0,1,\ldots,m=1,\ldots,M,$ (5)
$\quad\quad\quad\sum_{m=1}^{M}\sum_{j=1}^{k_{m}}\pi_{m,j}(t)n_{m,j}\leq N,~~t=0,1,2,\ldots,$ (6)
where $w_{m,j}\geq 0$ is the weight (e.g., priority) associated with the $(m,j)$-th inference task, and the discount factor $0<\gamma<1$ quantifies the diminishing importance of an inference task over time. At most $C_{m}$ feature generators for source $m$ can compute features at time $t$. Transmitting features for the $(m,j)$-th inference task requires $n_{m,j}$ of the $N$ wireless channels available. For each task, its inference error, $p_{m,j}(\Delta_{m,j}(t))$, depends on its AoI $\Delta_{m,j}(t)$ at time slot $t$, indicating the freshness of the feature used for inference.
@@ III-B Weakly Coupled MDP Formulation
The problem (4)-(6) is a weakly coupled Markov decision process (MDP) [13, 45, 47] with $K$ sub-MDPs (referred to as arms in the bandit literature), one per inference task $(m,j)$ across sources $m$. The state of each $(m,j)$-th MDP at each time $t$ is represented by the AoI $\Delta_{m,j}(t)$. The action is $\pi_{m,j}(t)$, and its per-timeslot cost is $\gamma^{t}p_{m,j}(\Delta_{m,j}(t))$ with discount factor $0<\gamma<1$. We can see that the state evolution defined in (3) and the cost of each MDP depends only on its current state and action. However, the actions for all MDPs $(\pi_{m,j}(t))_{\forall m,j}$ need to satisfy the constraints in (5)-(6). This interdependence of actions across MDPs through multiple resource constraints, despite independent state transitions and costs, makes the overall problem (4)-(6) a weakly coupled MDP [13, 45, 47]. Weakly coupled MDPs are PSPACE-hard because the number of states and actions grow exponentially with the number of sub-MDPs.
The restless multi-armed bandit (RMAB) problem is a special case of the weakly coupled MDP in (4)-(6). RMAB considers a single resource constraint, whereas our problem involves multiple resource constraints (5)-(6). The PSPACE-hard complexity of RMABs can be overcome by using Whittle indices [43, 10, 4, 32, 14, kadota2018scheduling, 23], gain indices [11, 36, 42], and linear programming-based indices [48] to construct asymptotically optimal policies, provided indexability and/or global attractor conditions are satisfied. However, these RMAB policies cannot be directly applied to our more general problem due to the presence of multiple resource constraints, which requires us to develop a new solution approach.
Solution approach. In Sec. IV&V, we follow the approach depicted in Fig. 2 to solve (4)-(6). We begin by deriving a relaxed Lagrangian problem. We then utilize the resulting solution to construct a maximum gain first (MGF) policy (Algorithm 1) for (4)-(6). Theorem 1 will demonstrate that the MGF policy becomes asymptotically optimal as the number of inference task $k_{m}$ for each source $m$ increases.